ADNAN MUHAMMED

Grade 12,

what is diffarence between Dot and Crossproducts ? and give equations for both Cross and Dotproducts

what is diffarence between Dot and Crossproducts? and give equations for both Cross and Dotproducts

Grade:11

2 Answers

sri tanish
8 years ago
Vector multiplication is of two types  – 1 } scalar multiplication{dot product }  2} vector multiplication – {cross product }

In Euclidean space, a Euclidean vector is a geometrical object that possesses both a magnitude and a direction. A vector can be pictured as an arrow. Its magnitude is its length, and its direction is the direction that the arrow points. The magnitude of a vector A is denoted by  \left\| \mathbf{A} \right\| . The dot product of two Euclidean vectors A and B is defined by[2][3]

\mathbf A \cdot \mathbf B = \left\| \mathbf A \right\| \, \left\| \mathbf B \right\| \cos \theta ,

where θ is the angle between A and B.  so basically dot product is a scalar . depending on cosine of angle b\w the vectors .  dot product is commutative .

 

CROSS PRODUCT:-

 

The magnitude of the cross product can be interpreted as the positive area of the parallelogram having a and b as sides (see Figure 1):

 
A = \left\| \mathbf{a} \times \mathbf{b} \right\| = \left\| \mathbf{a} \right\| \left\| \mathbf{b} \right\| \sin \theta. \,\!

SO CROSS PRODUCT  is a vector depending on sine of angle b\w the vectors . CROSS PRODUCT IS NOT COMMUTATIVE , RATHER IT IS ANTI COMMUTATIVE .

 

 

 

 

SHAIK HAFEEZUL KAREEM
8 years ago
Vector multiplication is of two types  – 1 } scalar multiplication{dot product }  2} vector multiplication – {cross product }

In Euclidean space, a Euclidean vector is a geometrical object that possesses both a magnitude and a direction. A vector can be pictured as an arrow. Its magnitude is its length, and its direction is the direction that the arrow points. The magnitude of a vector A is denoted by  \left\| \mathbf{A} \right\| . The dot product of two Euclidean vectors A and B is defined by[2][3]

\mathbf A \cdot \mathbf B = \left\| \mathbf A \right\| \, \left\| \mathbf B \right\| \cos \theta ,

where θ is the angle between A and B.  so basically dot product is a scalar . depending on cosine of angle b\w the vectors .  dot product is commutative .

 

CROSS PRODUCT:-

 

The magnitude of the cross product can be interpreted as the positive area of the parallelogram having a and b as sides (see Figure 1):

 
 

SO CROSS PRODUCT  is a vector depending on sine of angle b\w the vectors . CROSS PRODUCT IS NOT COMMUTATIVE , RATHER IT IS ANTI COMMUTATIVE .